# Imaginaries, products and the adele ring

### v2

@@ 1.1. Products and Boolean-valued models
Many thanks are due to Mariana Vicaria for her reading and comments on an earlier version of this paper, and to George Peterzil and Kobi Peterzil for their comments on this one.
@@ 1.2. Imaginaries
An imaginary sort of a first-order structure $M$ is a possible universe of a structure interpretable in $M$. The significance of imaginary sorts was established in the abstract setting of Classification Theory by Shelah [24]; subsequently all notions of stability theory, especially algebraic closure, were interpreted “in $T^{eq}$", i.e. with imaginary sorts taken into account. They were first studied for specific theories by Poizat [22]; he showed notably that the theories of algebraically closed fields and of differentially closed fields eliminate imaginaries, i.e. each imaginary sort is in definable bijection with a definable set of $n$-tuples. Poizat also pointed out that elimination of imaginaries for algebraically closed fields encapsulates, modulo general principles, classical Galois theory for fields.
Other theories have imaginary sorts that differ qualitatively from definable sets; an example is the value group of a valued field such as ${\mathbb{Q}}_{p}$. In this case it is a question of finding a useful basis for the imaginary sorts, meaning every imaginary sort can be embedded into a product of the basis elements. In the case of ${\mathbb{Q}}_{p}$, these would be the geometric sorts $GL_{n}({\mathbb{Q}}_{p})/GL_{n}({\mathbb{Z}}_{p})$ of lattices in ${\mathbb{Q}}_{p}^{n}$.
@@
At this point we can give a formulation of our main theorems. Let $T$ be any theory in a language $L$; assume for this introduction that each sort $S$ has at least one 0-definable element in $T$ (assumption $\clubsuit$; see Remark 1.3 Let $I$ be an infinite index set, and let $(M_{i}:i\in I)$ be a family of models of $T$. Let $M^{\sharp}=\prod_{i\in I}M_{i}$ with the natural structure (explained precisely in Example 2.8), including a sort $B$ for the Boolean algebra ${\mathop{\operatorname{\rm P}}}(I)$. For an equivalence relation $E$ on a finite product $D$ of sorts of $T$, let $E^{\sharp}$ and $E^{\flat}$ be equivalence relations on $D(M)$ defined by: $xE^{\sharp}y$ iff $x(i)Ey(i)$ for all $i\in I$, and $xE^{\flat}y$ iff $x(i)Ey(i)$ for all but finitely many $i\in I$. Let $M^{\flat}$ be the expansion of $M$ by a predicate for the ideal $I_{fin}\subset B$ of finite subsets of $I$.
@@
These descriptive-set theoretic ideas will also give us an important model-theoretic orthogonality principle. Let $I_{fin}$ be the ideal of finite subsets of the Boolean algebra ${\mathop{\operatorname{\rm P}}}({\mathbb{N}})$. Consider a product of structures $M=\prod_{i\in{\mathbb{N}}}M_{i}$, and define an equivalence relation $E_{fin}$ using the ideal $I_{fin}$, namely $aE_{fin}b$ iff $\{i:a_{i}=b_{i}\}\in I_{fin}$. In case $M_{i}$ is countable (but with at least two elements), $E_{fin}$ is a version of the paradigmatic non-smooth Borel equivalence relation $E_{0}$; Harrington, Kechris and Louveau show in [10] that a Borel equivalence relation is non-smooth if and only if $E_{0}$ can be reduced to it. In the adelic case, where $M_{i}$ is a local field, or more generally an uncountable Borel structure, $E_{fin}$ gives the equivalence relation known in descriptive set theory as $E_{1}$ (which is even less smooth.) On the other hand, we will see that the equivalence relations definable on a single $E_{fin}$-class are all smooth, in these cases. It will follow from this that there are no nontrivial definable maps or finitely multivalued maps from the reduced power $M/E_{fin}$ into $M$ (Proposition 5.3); this conclusion will hold in any model. This will allow us to study the imaginaries separately in each $E_{fin}$- class on various definable sets, equivalence relation, and in the quotient, and glue together the information.
@@
It would be very interesting to find a basis for the imaginaries for the algebraic integers, where a ’discrete’ field is also endowed from structure arising from embedding into an adele ring (see Problem 5.9.)
@@ 2. Preliminaries
Let L{L} be a language, possibly many-sorted. Let T{T} be a (possibly incomplete) theory of L{L}. It will be convenient to assume (and can easily be arranged by adding sorts) that any finite product of sorts stands in 0-definable bijection with a sort.
@@ 2.1. Language of $T^{boole}$
Let $L^{boole}$ be a language with the sorts of $L$ and one additional sort $B$. (For simplicity we will use notation as if $L$ has a single sort $K$.) On $B$ we have the operations of a Boolean algebra, and the equality relation. In addition, for any formula $\phi(x_{1},\ldots,x_{n})$ of $L$, we have a function symbol $\phi]:K^{n}\to B$ with the same variables. We view $\phi]$ as giving the Boolean truth value of $\phi$.
@@ 2.2. Axioms of $T^{boole}$
    If $T\models\forall x\phi$, then for all $a$, $\phi](a)=1$; where $x=(x_{1},..,x_{n})$ and $a=(a_{1},...,a_{n})$.
@@
    For any $a$, $\phi\mapsto[\phi](a)$ is a Boolean homomorphism. Also $[,]$ respects dummy variables, i.e. if $\phi(x,y):=\psi(x)$ then $\phi](a,b)=[\psi](b)$.
@@
    $(\forall b_{1},\ldots,b_{n})(\bigwedge_{i\neq j}b_{i}\cap b_{j}=0\&\bigwedge[\psi_{i}(y)]\geq b_{i}\implies(\exists x)\bigwedge_{i}[\phi_{i}(x,y)]\geq b_{i})$
@@
(4a) [Glueing] Whenever $b_{1},\ldots,b_{n}$ form a partition of unity in $B$, and $a_{1},\ldots,a_{n}$ are elements, there exists an element $a$ with $a=a_{i}]\geq b_{i}$.
(4b) If $T\models(\forall x)(\exists y)\phi(x,y)$ then $T^{boole}$ includes the axiom $(\forall x)(\exists y)\phi(x,y)]=1$.
It follows from these axioms that in fact $T^{boole}\models[\psi]=1$ whenever $T\models\psi$.
@@
Let $\alpha(x,u)$ be a formula of $T^{boole}$, where $x$ is a tuple of variables of sorts of $T$, and $u$ is a tuple of variables of sort $B$. Then for some formulas $\phi_{i}(x)$ of $T$ and $\beta$ of $B$,
$T^{boole}\models\alpha\iff\beta([\phi_{1}(x)],\ldots,[\phi_{k}(x)],u)).$
$T^{boole}$ is complete modulo the theory of $B$ in the language of Boolean algebras with additional constants, namely the elements $\psi]$ for sentences $\psi$ of $L$.
@@
Given $m^{\prime}\geq 1$, $L$-formulas $\phi^{\prime}_{1}(x),\dots,\phi^{\prime}_{m^{\prime}}(x)$ and a Boolean formula $\beta^{\prime}$ in $m^{\prime}$ variables, to construct $L$-formulas $\phi_{1}(x),\dots,\phi_{m}(x)$ (in the same free variables $x$) for some $m\geq 1$, and a Boolean formula $\beta$ in $m$ variables such that
@@
To see this, let $m=2^{m^{\prime}}$ and $s_{1},\dots,s_{m}$ be a list of subsets of $\{1,\dots,m^{\prime}\}$. For $k\leq m$, let
@@
For $l\leq m^{\prime}$, let $r_{l}:=\{k:k\leq m,~l\in s_{k}\}$, and define the Boolean formula $\beta:=\beta^{\prime}(\bigvee_{k\in r_{1}}u_{k},\dots,\bigvee_{k\in r_{m^{\prime}}}u_{k})$. Then $(\phi_{1},\dots,\phi_{m})$ is a partition of unity in $T$ and $\beta([\phi_{1}(x)],\dots,[\phi_{m}(x)])$ and $\beta^{\prime}([\phi^{\prime}_{1}(x)]\,\dots,[\phi^{\prime}_{m^{\prime}}(x)])$ are equivalent.
@@
To show this we may assume that $(\phi^{\prime}_{1},\dots,\phi^{\prime}_{m^{\prime}})$ is a partition of unity in $T$ by the above. Let $m=m^{\prime}$ and $\phi_{j}:=\exists x\phi^{\prime}_{j}(z)$ for $j\leq m$. Let $Part(u_{1},\dots,u_{m})$ express that $u_{1},\dots,u_{m}$ form a partition of unity in the Boolean algebra. Let
@@
where $\phi_{1},\ldots,\phi_{k}$ are a partition of unity in $T$, i.e. they are pairwise disjoint and $T\models(\forall x)(\bigvee_{i}\phi_{i})$. It is easy to see that $\Phi$ is closed under Boolean combinations. It follows immediately from the above argument that if $\phi(x,y)\in\Phi$ where $y$ is a single variable, then $(\exists y)\phi$ is $T^{boole}$-equivalent to a formula in $\Phi$. Quantifier-elimination for $T^{boole}$ follows. Stable embeddedness, and the statement on induced structure, follow from QE since there are no function symbols from $B$ into $M$. QE includes the case of zero variables, so that sentences of $T^{boole}$ are equivalent to sentences of $(B,[\psi])_{\psi}$ with $\psi$ varying over sentences of $L$.
@@ 2.4. Theories of Boolean algebras
Let $AB_{\forall}$ be the theory of Boolean algebras $B$ in a language with predicates $AT_{n},n=1,2,\cdots$. $AT_{1}$ is intended to denote the atoms, and $AT_{n}$ the sums of $n$ atoms. Let $ABA$ be the model completion. It is axiomatized by $AB_{\forall}$ along with axioms asserting that for any element $u$ of $AT_{n}$ there exist atoms $a_{1},\ldots,a_{n}$ with $u=\sum a_{i}$; and that for any $u>0$ there exists $a\in AT_{1}$ with $a<u$. Let $ABA_{n}$ assert that there are exactly $n$ atoms, while $ABA_{\infty}$ asserts there are infinitely many.
Let ABAI be the expansion of $AB_{\forall}$ obtained by expanding the Boolean algebra $B$ by adding a predicate $J$ for an ideal of $B$ containing all atoms, as well as a sort $\bar{B}=B/J$, and a function symbol for the natural map $B\to\bar{B}$ with kernel $J$. Let ABAI{{ABAI}} be ${{ABAI}}_{\forall}$ along with an additional axiom asserting that $B\to{\bar{B}}$ is surjective, and $\bar{B}$} is atomless. Let ${ABAI}_{\infty}$ assert, in addition, that there are infinitely many atoms of $B$. We define $ABAI_{n}=ABA_{n}+({\bar{B}}~\mathrm{is~atomless}.).$ We note that if $B/J$ is atomless, then $J$ contains all the atoms of $B$.
@@
    ${{ABAI}}_{\infty}$ and each $ABAI_{n}$ is complete and eliminates quantifiers.
@@
(1) is [7], Theorem 2.3, and presumably (2) is not hard to deduce. Alternatively (2) can be proved using the criterion of ([4], Appendix) that a 0-definable $D$ is embedded and stably embedded in $T$ if $T$ has a saturated model $M$ such that $Aut(M)\to Aut(D^{M})$ is surjective. (This will also gives an alternative proof of (1).)
@@
(4) Clear from completeness of $ABA_{\infty}$ and each $ABA_{n}$.

### v3

@@ 1.1. Products and Boolean-valued models
Many thanks to Mariana Vicaria for her extended and comments on an earlier version of this paper, to George Peterzil and to Kobi Peterzil, Mira Tartarotti, and to the anonymous referee; the paper was much improved by his or her suggestions.
@@ 1.2. Imaginaries
An imaginary sort of a first-order structure $M$ is a possible universe of a structure interpretable in $M$. It can be presented as a quotient $D/E$, with $D$ a definable set (perhaps of tuples) and $E$ a definable equivalence relation on $D$. The significance of imaginary sorts was established in the abstract setting of Classification Theory by Shelah [26]; subsequently all notions of stability theory, especially algebraic closure, were interpreted “in $T^{eq}$”, i.e. with imaginary sorts taken into account. They were first studied for specific theories by Poizat [23]; he showed notably that the theories of algebraically closed fields and of differentially closed fields eliminate imaginaries, i.e. each imaginary sort is in definable bijection with a definable set of $n$-tuples. Poizat also pointed out that elimination of imaginaries for algebraically closed fields encapsulates, modulo general principles, classical Galois theory for fields.
Other theories have imaginary sorts that differ qualitatively from definable sets; an example is the value group of a valued field such as ${\mathbb{Q}}_{p}$. In this case it is a question of finding a useful basis for the imaginary sorts, meaning every imaginary sort can be definably embedded into a product of the basis elements. In the case of ${\mathbb{Q}}_{p}$, these would be the geometric sorts $GL_{n}({\mathbb{Q}}_{p})/GL_{n}({\mathbb{Z}}_{p})$ of lattices in ${\mathbb{Q}}_{p}^{n}$.
@@
At this point we can give a formulation of our main theorems. Let $T$ be a first-order theory in a language $L$; $T$ may be incomplete. We assume for this introduction that each sort $S$ has at least one 0-definable element in $T$ (assumption $\clubsuit$; see Remark 1.3). Let $I$ be an infinite index set, and let $(M_{i}:i\in I)$ be a family of models of $T$. Let $M^{\sharp}=\prod_{i\in I}M_{i}$ with the natural structure (explained precisely in Example 2.9), including a sort $B$ for the Boolean algebra ${\mathop{\operatorname{\rm P}}}(I)$. For an equivalence relation $E$ on a finite product $D$ of sorts of $T$, let $E^{\sharp}$ and $E^{\flat}$ be equivalence relations on $D(M)$ defined by: $xE^{\sharp}y$ iff $x(i)Ey(i)$ for all $i\in I$, and $xE^{\flat}y$ iff $x(i)Ey(i)$ for all but finitely many $i\in I$. Let $M^{\flat}$ be the expansion of $M$ by a predicate for the ideal $I_{fin}\subset B$ of finite subsets of $I$.
@@
These descriptive-set theoretic ideas will also give us an important model-theoretic orthogonality principle. Let $I_{fin}$ be the ideal of finite subsets of $I$, and $E_{fin}$ the associated equivalence relation: $aE_{fin}b$ iff $\{i\in I:a_{i}\neq b_{i}\}\in I_{fin}$. In case $M_{i}$ is countable (but with at least two elements), $E_{fin}$ is a version of the paradigmatic non-smooth Borel equivalence relation $E_{0}$; Harrington, Kechris and Louveau show in [11] that a Borel equivalence relation is non-smooth if and only if $E_{0}$ can be reduced to it. In the adelic case, where $M_{i}$ is a local field, or more generally an uncountable Borel structure, $E_{fin}$ gives the equivalence relation known in descriptive set theory as $E_{1}$ (which is even less smooth.) On the other hand, we will see that the equivalence relations definable on a single $E_{fin}$-class are all smooth, in these cases. It will follow from this that there are no nontrivial definable maps or finitely multivalued maps from the reduced power $M/E_{fin}$ into $M$ (Proposition 5.3); this conclusion will hold in any model. This will allow us to study the imaginaries separately in each $E_{fin}$- class on various definable sets, equivalence relation, and in the quotient, and glue together the information.
@@
It would be very interesting to find a basis for the imaginaries for the algebraic integers, where a ‘discrete’ field is also endowed from structure arising from embedding into an adele ring (see Problem 5.10).
@@ 2. Preliminaries
Let L{L} be a language, possibly many-sorted.
Let T{T} be a first-order theory of L{L}. It will be convenient to assume (and can easily be arranged by adding sorts) that any finite product of sorts stands in 0-definable bijection with a sort. Let $L_{BA}=\{\cup,\cap,0,1,-\}$ be the language of Boolean algebras. We denote the Boolean complement of an element $a$ by 1a1-a.
Recall that in a given structure $M$, the algebraic closure of a subset $A$ is the union of all finite $M_{A}$-definable sets; while the definable closure is the set of all $M_{A}$-definable elements, equivalently the union of all $M_{A}$-definable one-element sets. We denote algebraic and definable closure by $\mathop{\rm acl}\nolimits$ and $\mathop{\rm dcl}\nolimits$, respectively. While our goal is to eliminate abstract imaginary sorts in favour of concrete ones, as noted in § 1.2, it will sometimes be useful to consider (abstract) imaginary elements; the notions of $\mathop{\rm acl}\nolimits,\mathop{\rm dcl}\nolimits$ apply also to $M^{eq}$. Note that the relation $a\in\mathop{\rm acl}\nolimits(b)$ between given elements is unambiguous - in case $a,b$ are ‘real’, the relation holds in $M$ if and only if it holds in $M^{eq}$, and likewise for $\mathop{\rm dcl}\nolimits$. In one or two arguments it is essential to take the closure in $M$ - restrict to ‘real’ elements; we will then explicitly say so.
@@ 2.1. Language of $T^{boole}$
Let $L^{boole}$ be a language with the sorts of $L$ and one additional sort $B$ equipped with the language $L_{BA}$. (For simplicity we will use notation as if $L$ has a single sort $K$.) On $B$ we have the operations of a Boolean algebra, and the equality relation. In addition, for any formula $\phi(x_{1},\ldots,x_{n})$ of $L$, we have a function symbol $\phi]:K^{n}\to B$ with the same variables. We view $\phi]$ as giving the Boolean truth value of $\phi$. In (2-4) below, all variables $x,y,a,b,\ldots$ range over the sorts of $L$.
@@ 2.2. Axioms of $T^{boole}$
    If $T\models\forall x\phi$, then for all $a$ from $K$, $\phi](a)=1$; where $x=(x_{1},..,x_{n})$ and $a=(a_{1},...,a_{n})$.
@@
    For any tuple $a$ from $K$, $\phi\mapsto[\phi](a)$ is a Boolean homomorphism. Also $[,]$ respects dummy variables, i.e. if $\phi(x,y):=\psi(x)$ then $\phi](a,b)=[\psi](b)$. (We will sometimes denote $\phi](z)$ by $\phi(z)]$ when this simplifies the notation.)
@@
    $(\forall b_{1},\ldots,b_{n})(\bigwedge_{i\neq j}b_{i}\cap b_{j}=0\wedge\bigwedge_{i\leq n}[\psi_{i}(y)]\geq b_{i}\implies(\exists x)\bigwedge_{i\leq n}[\phi_{i}(x,y)]\geq b_{i})$
@@
(4a) [Glueing] Whenever $b_{1},\ldots,b_{n}$ form a partition of unity in $B$, and $a_{1},\ldots,a_{n}\in K$ are elements, there exists an element $a$ with $a=a_{i}]\geq b_{i}$.
(4b) If $T\models(\forall x)(\exists y)\phi(x,y)$ then $T^{boole}$ includes the axiom $(\forall x)(\exists y)[\phi(x,y)]=1$.
It follows from these axioms that in fact $T^{boole}\models[\psi]=1$ whenever $T\models\psi$ (for $\psi$ an $L$-sentence).
We call a sequence of $L$-formulas $\phi_{1}(x),\dots,\phi_{k}(x)$ a partition of unity in $T$ if for $i\neq j$, $T\models\neg(\exists x)(\phi_{i}\wedge\phi_{j})$, and $T\models(\forall x)(\bigvee_{i\leq k}\phi_{i}(x))$
@@
Let $\alpha(x,u)$ be a formula of $T^{boole}$, where $x$ is a tuple of variables of sorts of $T$, and $u$ is a tuple of variables of sort $B$. Then for some $L$-formulas $\phi_{1}(x),\dots,\phi_{k}(x)$, that form a partition of unity in $T$, and some $L_{BA}$-formula $\beta$,
$T^{boole}\models\alpha\iff\beta([\phi_{1}(x)],\ldots,[\phi_{k}(x)],u).$
$T^{boole}$ is complete modulo the theory of $B$ in the language of Boolean algebras $L_{BA}$ with additional constants, namely the elements $\psi]$ for sentences $\psi$ of $L$.
@@
We start with
Claim .  Given $m^{\prime}\geq 1$, $L$-formulas $\phi^{\prime}_{1}(x),\dots,\phi^{\prime}_{m^{\prime}}(x)$ and a Boolean formula $\beta^{\prime}$ in $m^{\prime}$ variables, there are $L$-formulas $\phi_{1}(x),\dots,\phi_{m}(x)$ (in the same free variables $x$) for some $m\geq 1$, and a Boolean formula $\beta$ in $m$ variables such that
@@
Proof.
Let $m=2^{m^{\prime}}$ and $s_{1},\dots,s_{m}$ be a list of subsets of $\{1,\dots,m^{\prime}\}$. For $k\leq m$, let
@@
For $l\leq m^{\prime}$, let $r_{l}:=\{k:k\leq m,~l\in s_{k}\}$, and define the Boolean formula $\beta:=\beta^{\prime}(\bigvee_{k\in r_{1}}u_{k},\dots,\bigvee_{k\in r_{m^{\prime}}}u_{k})$. Then $(\phi_{1},\dots,\phi_{m})$ is a partition of unity in $T$ and $\beta([\phi_{1}(x)],\dots,[\phi_{m}(x)])$ and $\beta^{\prime}([\phi^{\prime}_{1}(x)]\,\dots,[\phi^{\prime}_{m^{\prime}}(x)])$ are equivalent. This proves the Claim.
@@
To show this we may assume that $(\phi^{\prime}_{1},\dots,\phi^{\prime}_{m^{\prime}})$ is a partition of unity in $T$ by the above. Let $m=m^{\prime}$ and $\phi_{j}:=\exists x\phi^{\prime}_{j}$ for $j\leq m$. Let $Part(u_{1},\dots,u_{m})$ express that $u_{1},\dots,u_{m}$ form a partition of unity in the Boolean algebra. Let
@@
where $\phi_{1}(x),\ldots,\phi_{k}(x)$ are a partition of unity in $T$. It is easy to see that $\Phi$ is closed under Boolean combinations. It follows immediately from the above argument that if $\phi(x,y)\in\Phi$ where $y$ is a single variable, then $(\exists y)\phi$ is $T^{boole}$-equivalent to a formula in $\Phi$. Quantifier-elimination for $T^{boole}$ follows. Stable embeddedness, and the statement on induced structure, follow from QE since there are no function symbols from $B$ into $\mathcal{M}$. QE includes the case of zero variables, so that sentences of $T^{boole}$ are equivalent to sentences of $(B,[\psi])_{\psi}$ with $\psi$ varying over sentences of $L$.
@@ 2.4. Theories of Boolean algebras
Let $AB_{\forall}$ be the theory of Boolean algebras $B$ in the language of Boolean algebras $L_{BA}$ augmented by predicates $AT_{n},n=1,2,\cdots$; $AT_{1}$ is intended to denote the atoms, and $AT_{n}$ the sums of $n$ atoms, so this is merely an extension by definition. Let $ABA$ be the model completion. It is axiomatized by $AB_{\forall}$ along with axioms asserting that for any element $u$ of $AT_{n}$ there exist atoms $a_{1},\ldots,a_{n}$ with $u=\sum_{i\leq n}a_{i}$; and that for any $u>0$ there exists $a\in AT_{1}$ with $a\leq u$. Let $ABA_{n}$ assert that there are exactly $n$ atoms, while $ABA_{\infty}$ asserts there are infinitely many. We will denote the theory of atomless Boolean algebras in the language of Boolean algebras by $BA_{atomless}^{boole}$.
Let ABAI be the expansion of $AB_{\forall}$ obtained by expanding the Boolean algebra $B$ by adding a predicate $J$ for an ideal of $B$ containing all atoms, as well as a sort $\bar{B}=B/J$, and a function symbol for the natural map $B\to\bar{B}$ with kernel $J$. Let ABAI{{ABAI}} be ${{ABAI}}_{\forall}$ along with an additional axiom asserting that $B\to{\bar{B}}$ is surjective, and $\bar{B}$} is atomless. Let ${ABAI}_{\infty}$ assert, in addition, that there are infinitely many atoms of $B$. We define $ABAI_{n}=ABA_{n}+({\bar{B}}~\mathrm{is~atomless}).$ We note that if $B/J$ is atomless, then $J$ contains all the atoms of $B$.
@@
    $BA_{atomless}^{boole}$, ${{ABAI}}_{\infty}$ and each $ABAI_{n}$ is complete and eliminates quantifiers.
@@
(1) These results are well-known. For a proof for the case of $BA_{atomless}^{boole}$ see [24, Theorem 6.22, p.85]. For a proof for the theories $BA_{atomless}^{boole}$, ${{ABAI}}_{\infty}$ and each $ABAI_{n}$ see [8].
(2) can be proved directly from (1) or by using the criterion of Lemma 1 in the Appendix of [5]: an $\emptyset$-definable $D$ is embedded and stably embedded in $T$ if $T$ has a saturated model $M$ such that $Aut(M)\to Aut(D^{M})$ is surjective. This also gives an alternative proof of (1).
@@
(4) This follows from completeness of $ABA_{\infty}$.